OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (espex):

Work this just like any other algebra problem, get F by itself then plug in your 20 degrees and solve.

OpenStudy (anonymous):

idk how to do that tho,

OpenStudy (espex):

\[C=\frac{5}{9}(F-32)\] So you start by dividing both sides by 5/9, which is really just multiplying by 9/5.\[\frac{9}{5}C=F-32\] Next you add 32 to both sides, leaving F by itself.\[\frac{9}{5}C+32=F\]